A vacuum gauge reads 60.2 cm Hg. Determine the absolute pressure in kPa if the barometer reads 75.7 cm Hg.
1
Expert's answer
2021-10-28T12:34:39-0400

"P_{vacuum \; gauge} = 60.2 \;cmHg \\

P_{atm} = 75.7 \;cmHg \\

P_{absolute} = -P_{vacuum \; gauge} + P_{atm} \\

= -60.2 + 75.7 \\

= 15.5 \; cmHg"

We know that 1 cm Hg = 1333.2 Pa

"P_{absolute} = 15.5 \times 1333.2 = 20664.6 \;Pa \\

P_{absolute} = 20.664 \;kPa"